OpenStudy (anonymous):

Factor: x2 - 16 (x - 4)(x - 4) (x - 4)(x + 4) (x - 8)(x + 2) (x - 2)(x + 8)

OpenStudy (lgbasallote):

it's a difference of two squares..what is the square root of x^2 and 16 (individually)?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

and 4

OpenStudy (unklerhaukus):

\[(x - 4)(x - 4)=x^2-4x-4x+16\]\[(x - 4)(x + 4)=x^2-4x+4x-16\]\[(x - 8)(x + 2)=x^2+2x-8x-16\]\[(x - 2)(x + 8)=x^2+8x-2x-16\]

OpenStudy (lgbasallote):

^that's a fine way too....

OpenStudy (anonymous):

I have no idea what that meant... but ok

OpenStudy (lgbasallote):

it meant trial and error =))

OpenStudy (anonymous):

Don't know what that means.

OpenStudy (lgbasallote):

you solve each of the choices and see

OpenStudy (unklerhaukus):

i have just done the expansion of each expression

OpenStudy (anonymous):

(x - 4)(x + 4)